A normally distributed random variable X has a mean of 120 and a standard deviation of 8. Find the standard deviation normal values of the following scores:


a.200


b.70


c.150


d.300

1
Expert's answer
2022-03-08T01:49:13-0500
"Z=\dfrac{X-\mu}{\sigma}\sim N(0,1)"

Given "\mu=120, \sigma=8"

a.


"Z=\dfrac{200-120}{8}=10"

b.


"Z=\dfrac{70-120}{8}=-6.25"

c.


"Z=\dfrac{150-120}{8}=3.75"

d.


"Z=\dfrac{300-120}{8}=22.5"